# An Extensive Study on D2C: Overfitting Remediation in Deep Learning Using a Decentralized Approach

### v1

@@ IV-B Training & Averaging of Trained Parameters
  Input: Training dataset $D$, Number of subsets $N$, Local epochs $E$, Global epochs $E_{global}$, Batch size $B$, Learning rate $lr$
  Output: Central model $M_{c}$ with aggregated weights
  function Divide2Conquer($D$, $N$, $E$, $E_{global}$, $B$, $lr$)
        Shuffle the dataset $D$
@@
        Divide $D$ into $N$ subsets: $\{D_{1},D_{2},\dots,D_{N}\}$, maintaining class distributions
        Initialize central model $M_{c}$ with parameter vector $\theta_{c}$
  for $i=1$ to $N$ do
           Initialize edge model $M_{i}$ with $\theta_{i}=\theta_{c}$
           Set scaling factor $s_{i}=\frac{|D_{i}|}{|D|}$
  end for
  for $g=1$ to $E_{global}$ do
           Initialize weight accumulator $W_{central}=0$
     for $i=1$ to $N$ do
                  Set $\theta_{i}=\theta_{c}$
        for $e=1$ to $E$ do
                         Train $M_{i}$ on $D_{i}$ using batch size $B$ and learning rate $lr$
        end for
@@
                  Obtain updated weights $\theta_{i}$ from $M_{i}$
                  Scale weights: $\theta_{i}=s_{i}\cdot\theta_{i}$
                  Accumulate scaled weights into $W_{central}$: $W_{central}=W_{central}+\theta_{i}$
     end for
           Compute averaged central weights:
$\theta_{c}=\frac{1}{\sum_{i=1}^{N}s_{i}}W_{central}$
           Update central model $M_{c}$ with new weights $\theta_{c}$
  end for
  return central model $M_{c}$
  end function
IV-C Global Hyperparameter Tuning
Tuning the overall model consists of two stages. Firstly, the edge model and its hyperparameters should be tuned using a subset/entire training set and the validation set to find a suitable model to train each training subset. Next, that base model is utilized to tune the Central model. The new global hyperparameters that need tuning, in this case, are the Number of Subsets of the Training Set and the Number of Epochs before each round of Global Averaging. Testing for different values for these hyperparameters, the implementation, and the evaluation of our method are done using the best combination. The appropriate number of training subsets can be determined first by varying the number of training subsets and observing the performance metrics like test/validation accuracy, F1 score, log loss, and also the validation loss curve. The appropriate number of training epochs in each subset before the averaging is done each time is likely to be a small number $(1\sim 3)$ since the edge model can quickly overfit the training subset with a reduced size. So, it makes more sense to set a small number of epochs initially and tune the number of training subsets. Once the number of subsets is determined, the same process can be repeated by varying the number of training epochs and determining the appropriate value of it.
Figure 3: D2C at a Glance – Training on each subset, averaging the trained parameters, setting the Central model parameters to these values, and sharing the averaged parameters back to the Edge models to keep the loop running.
V Experimental Setup
V-A Datasets
@@
For image classification tasks, we employed three widely used benchmark datasets: CIFAR-10 [14], MNIST [7], and Fashion MNIST [12]. CIFAR-10 provides a challenging image classification task with 60,000 32x32 RGB images across 10 categories. MNIST offers a simpler task with 70,000 28x28 grayscale images of handwritten digits. Fashion MNIST presents a more complex variation with 70,000 28x28 grayscale images of fashion items. To address the overfitting challenge in facial expression recognition, we utilized the FER-2013 [16] dataset, which comprises over 35,000 grayscale images of seven basic emotions. The diverse and noisy nature of this internet-sourced dataset makes it particularly suitable for evaluating generalization performance.
For audio-based tasks, we created a comprehensive dataset by combining TESS [19], CREMA-D [20], and RAVDESS [21]. This dataset encompasses 14 emotion classes, offering a diverse range of emotional expressions and recording conditions. By merging these datasets, we aim to improve model generalization by mitigating biases and limitations inherent in individual datasets.
To test our method on text data, we employed the AG News [22] dataset, which consists of approximately 127,600 news articles categorized into four balanced classes: World, Sports, Business, and Science/Technology. This dataset provides a large-scale, real-world benchmark for evaluating text classification models, particularly in the context of Big Data.
V-B Experimental Details
@@
1. Image Classification: At first, all the images are reshaped and normalized. We applied one-hot encoding to the labels to use cross-entropy loss later. We split the training data into training and validation sets using Scikit-Learn. We then divided the training set into multiple subsets. Before constructing the model, processing each training subset into tensors, and batching them was our last step. Each of our edge models was comprised of 4 Convolutional (32 to 256 kernels progressively) and MaxPooling layer (2X2) blocks followed by one fully connected hidden layer (128 neurons). Each block contained two identical convolutional layers and one pooling layer.
@@
In all cases, we used a 90-10 Training–Validation split, Adam optimizer, and applied dropouts and batch normalization layers. This is important to test for D2C’s effectiveness when used on top of other methods that address overfitting. We used the predefined test sets for testing.
VI Results and Discussions
We will use loss and accuracy curves, Accuracy, F1 Score, AUC-ROC, and log loss to analyze the results and evaluate our method. We also varied the two new global hyperparameters we mentioned in the previous section: the Number of Subsets of the Training Set and the Number of Epochs before each round of Central Averaging. To refer to them concisely, we will use the variables $\boldsymbol{N}$ and $\boldsymbol{E}$ respectively in this section.
@@ VI Results and Discussions
Figure 4: Decision boundary Visualization with the Traditional approach vs. the Divide2Conquer method.
Figure 5: Decision Boundary from the edge models trained on each subset after dividing the training data into 3 subsets applying the D2C method.
VI-A Visualizing Decision Boundary
At first, we will visualize the change in decision boundaries due to applying the D2C method and analyze the findings. We used a synthetic binary classification dataset to simulate a simple decision boundary. This dataset was created using the Scikit-Learn library. Then the 240-sample large dataset was divided into a training and a test set using a 75-25 split. The neural network architecture we used for these experiments comprised three hidden layers having 100 neurons each. No regularization or data augmentation was used for decision boundary visualization, to observe the effect of D2C explicitly.
At first, we adopted the centralized, traditional approach and fed the training dataset to the sole model. After training, we got the decision boundary shown in Figure 4a.
As we can see, the decision boundary for the traditional approach is highly complex and non-linear here. The model creates intricate regions to distinguish between the two classes. The boundary closely follows individual points, resulting in a jagged, convoluted shape. The complexity of the decision boundary suggests that the model is overfitting to the training data. It is trying too hard to separate every point, including noise, rather than focusing on general patterns in the data. Evidence of overfitting is seen in how the model creates small pockets of blue in the red region and vice versa. This indicates the model has likely memorized the data rather than learning generalizable patterns. This means that this approach is likely to produce excellent results on training data. Still, it is likely to struggle with new, unseen data due to its sensitivity to specific details in the training set. The highly irregular boundary reflects poor generalization ability.
Then, to see the effect of D2C on decision boundaries, we divided our training set into multiple subsets and trained them parallelly using identical Edge Models. Figure 5 shows the decision boundaries we got from each edge model.
@@ VI-A Visualizing Decision Boundary
To verify, the trained weights from each subset were averaged to get the final weights for the Central Model. The resulting decision boundary is shown in Figure 4b.
We can see that the decision boundary from the central model of D2C is much simpler, dividing the feature space almost diagonally. This suggests a less flexible, more generalized model due to averaging the weights from multiple subsets. This boundary is smooth and does not follow the exact positions of individual data points as closely. The model is more generalized, focusing on the overall trend in the data rather than trying to accommodate every single datapoint, exhibiting much less overfitting compared to what we saw in the traditional approach using the same model architecture. By averaging the weights from different subsets, the model smooths out the noise in the data and prevents overfitting to specific training samples. An individual noisy sample does not affect all the edge models in D2C, and its impact is diluted through averaging. This leads the central model to create a decision boundary based on the more common patterns present in all the subsets, ensuring better generalization. Further proof of this generalization is seen in the training and test accuracies for the traditional approach and D2C using different numbers of subsets. We summarize these accuracies in Table I.
TABLE I: Performance of the traditional approach vs different values of $N$ Using the D2C method.
@@
Our observations prove that models using the D2C approach perform better on unseen data, as they avoid overfitting by keeping the decision boundary simple and focusing on capturing the broader, more consistent patterns in the data. So, D2C can be seen as yet another technique to reduce overfitting. However, several other techniques address overfitting. So, whether using D2C on top of other regularization and augmentation techniques (such as dropouts) improves the performance of a model is the key question in application level. In the following subsections, we will investigate this using the architectures (incorporating regularization methods like dropouts in the edge models) we described in the previous section on several datasets from different domains.
VI-B Analyzing Loss & Accuracy Curves
@@
We will analyze the results obtained from our experiments from the 3 different domains. The loss curve is a definitive indicator of overfitting. As we saw in Eq. (6), The validation/test loss truly represents the true error of the model. So we will monitor the validation loss curves from the different experiments we performed to compare and analyze if D2C improves the generalizing performance, in the form of minimizing the validation loss as the training progresses.
@@
Image Classification: Let us first take a look at the loss curves generated using the CIFAR-10, Fashion MNIST, MNIST, and FER2013 datasets. For now, we will keep the number of epochs before each round of central averaging, $\boldsymbol{E}$, equal to $1$. It will ensure a proper comparison between the traditional method and our proposed method while varying the number of subsets, $\boldsymbol{N}$. In our experiments, how much we varied $N$ depended on the performance (Loss curve, Accuracy) trend as we increased $N$, the number of subsets. In the case of all the datasets, we stopped varying $N$ whenever we observed a definitive performance drop with increasing N. For example, in the case of CIFAR-10, we varied $N$ in the range of $1$ to $9$.
@@
Figure 6: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on CIFAR-10.
@@
Figure 7: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on Fashion MNIST.
@@
Figure 8: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on MNIST.
@@
Figure 9: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on FER2013.
@@
As we can see from Figure 6 and Figure 9, the validation loss keeps decreasing with each epoch for longer when using our proposed D2C method on these datasets but starts to increase after very few epochs again in the case of traditional NN implementations using the entire training set. In Figure 7, we can see that D2C with 9 subsets almost entirely eliminates the trend of an increasing loss after a few epochs. In machine learning, overfitting is marked by a model’s increasing validation loss after a certain number of training epochs, indicating that the model has started to capture noise and specific variations in the training set that do not generalize to new data. A model that reaches this overfitting point later, meaning it maintains a lower validation loss over a longer period, suggests it is better regularized or has a complexity that is well-suited to the underlying patterns in the data. A model with a delayed overfitting point continues to find generalizable patterns within the data through many epochs, highlighting that it remains in a generalization phase for longer, rather than immediately entering an overfitting phase. We can also see that, in all these cases, dividing the training data into more and more subsets delays and slows down the abrupt increase in validation loss as the training progresses. For example, in the case of CIFAR-10, the loss curve generated using the traditional approach without dividing the training data (no subsets) reaches its minimum, $0.536$, at around $16$ epochs. Then it keeps increasing and reaches $0.77$ by the end of 100 epochs. As we set the number of epochs before each round of central averaging, $E=1$, each central epoch is equivalent to one epoch in the traditional approach. For the number of subsets, $N=3$, the loss curve reaches its minimum, $0.462$, at around $21$ epochs and then it keeps increasing and reaches0.640.640.640.64 by the end of 100 epochs. When we further divided the training data into more subsets, setting $N=9$, the loss curve reaches its minimum,$0.47$, at around 46 epochs and then it keeps increasing and reaches $0.585$ by the end of 100 epochs. So, clearly, the convergence becomes slower, and the abrupt increase of validation loss is minimized as we divide the training set into more and more subsets, which indicates reduced overfitting.
Also, D2C manages to maintain a lower validation loss in general in all cases, including MNIST as we can see from Figure 8. The minimum value of loss that each of the cases discussed above also tells a similar story. The more subsets we divide the training set into, the less the validation loss we can observe from the curves. When a model maintains a lower validation loss during training, it indicates that the model is able to generalize well to unseen data. Theoretically, overfitting occurs when a model becomes overly complex relative to the data it is learning, allowing it to capture not only meaningful patterns but also specific noise or minor irregularities unique to the training set. This focus on noise causes the model to perform poorly on new data, which is reflected in a higher validation loss or an accelerated rise in validation loss. In contrast, a model that keeps validation loss lower while training likely has a structure that is well-suited to the data’s true underlying patterns without becoming excessively sensitive to its unique characteristics. By maintaining a lower validation loss, the model demonstrates an ability to generalize—learning patterns that extend beyond the training set and apply well to validation data. This suggests that the model has achieved a balance, where it is complex enough to capture essential relationships within the data but not so complex that it becomes reliant on details that do not generalize. Consequently, a model that sustains a lower validation loss is theoretically overfitting less, as it shows a stronger alignment with the data’s general structure rather than its idiosyncrasies. So, we can conclude that our proposed method manages to bring down the validation loss during training compared to the traditional method, which is a definitive sign of better generalization and reduced overfitting. We can further conclude that this reduction in validation loss becomes more apparent as we increase $N$, the number of subsets. The most clear illustration of this phenomenon is seen in the case of FER2013 (Figure 9). The Facial Expression Recognition task using the FER2013 dataset is prone to overfitting to a great extent. Dividing the training set into an increasing number of subsets evidently minimizes that overfitting. So, dividing the training set into more and more subsets helps reduce overfitting, an observation that is consistent with what we saw in Eq. (13). The more the number of subsets ($k$ in that equation), the less the variance, s/k, for the best-case scenario. Hence, with the decrease in variance comes the decrease in overfitting. Also, as we saw before, the true error, represented by validation loss in this case, is minimized using D2C, which proves our hypothesis and shows that D2C is indeed effective in treating the problem of overfitting.
Figure 10: Validation accuracy curves for the traditional method & the D2C method for different numbers of subsets, N, on Fashion MNIST.
However, the loss curve isn’t the only important factor during the training process. The relation between log loss and accuracy may not be so linear. That is why it is important to observe the validation accuracy too. As we can see from Figure 10, for all the different values of the number of subsets, $N$, even though the loss reaches its minimum much earlier, the accuracy shows an increasing trend (even if slightly) almost throughout the entire training process, right up until the 60th epoch. This shows why continuing training for a significant period might be necessary even if the validation loss starts increasing. This also shows why early stopping based on loss curves may not be a good idea in both cases: the traditional approach and the D2C method. This is something we observed in the cases of all these datasets. Also, we can notice a visible improvement in terms of validation accuracy in Figure 10. At the very least, dividing the training set into multiple subsets ensures that the maximum accuracy is attained while keeping the log loss relatively in control, which characterizes a more confident and robust model with better generalization ability.
@@
There is another interesting aspect of the validation loss curve here. As we can see from Figure 8, the validation loss curve generated using the traditional approach deviates much more than the ones generated using the D2C method. A curve that does not fluctuate much after converging suggests that the model has reached a stable state. This stability is often a sign of robustness, meaning the model is not overly sensitive to slight variations in the data or training process. Dividing the dataset into more and more subsets results in smaller and smaller deviations in the validation loss curve. This was evident in previous loss curves too, to a lesser extent. This proves that dividing the training set into multiple subsets not only improves accuracy and loss but also brings stability and consistency to the loss/accuracy curves.
Let us also take a look at the validation loss curves generated using FER2013 while keeping the number of subsets, $N$, constant and varying the number of local epochs before each round of central averaging, $E$. We varied $E$ between $1$ to $3$ and Kept $N$ constant at $2$ and $3$ separately.
Figure 11: Validation loss curves for different numbers of epochs before each round of central averaging, while keeping the number of subsets, N = 2 & 3.
We can see from Figure 11 that when we vary the number of epochs before each round of central averaging, $E$, there is no clear indication of any kind of change in the trend or the rate of change in the validation losses as the training progresses. This is unlike when we varied the number of subsets, $N$ keeping $E$ constant, where we saw that increasing the number of subsets delayed and minimized the increase in validation loss as the training progressed quite clearly. Also, the convergence got slower. But here, for both $N=2$ (shades of blue) and $N=3$ (shades of green), the trends from the validation curves are quite similar for different values of $E$ throughout the training phase. In light of the above observations, we can conclude that varying the number of epochs before each round of central averaging, $E$ does not have any obvious implications on log loss as we saw in the case of varying the number of subsets, $N$. However, an increased number of local epochs before each round of central averaging means more scope for learning from each of the subsets before averaging, sometimes even overfitting on a particular subset a bit more. We will observe the effect of varying the number of epochs further in the later subsections when we observe the final accuracy and log loss in each case. The implication of varying $E$ is likely to be different on different datasets from different domains as we will see in the case of the number of subsets too.
Audio Classification: Overfitting is a common problem in audio classification tasks. But it is more apparent when the model is trained on a particular dataset, and then tested on data from a different source. One way of tackling this issue is by combining datasets to increase the diversity in both training and test/validation data. We combined the datasets TESS [19], CREMA-D [20], and RAVDESS [21] for this task. Let us take a look at the validation loss and accuracy curves generated using this Combined Audio dataset.
Figure 12: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on the Combined Audio dataset.
As we can see from Figure 12, there is overfitting to some extent in all cases, which causes the validation loss to creep up after a few epochs but the rate at which this phenomenon takes place is much less when the training set is divided into multiple subsets. Once again, the rate of increase in validation loss goes down as we increase the number of subsets. Even though we converted the 1D time series audio files to 2D Log-Mel spectrograms to form something like an image, the classification based on these Mel-specs differs greatly from traditional image classification. So, the overfitting-resistant nature of models using our proposed method being apparent in this case too is a very promising sign. However, the small size of this dataset may not be ideal in the context of applying D2C. We will analyze it in detail in the next subsection based on several performance metrics.
Text Classification: Text classification is an important case study when it comes to testing the Divide2Conquer method. For both image classification and audio classification, we used 2D data and CNN to carry out the tasks. But text classification deals with 1D sequences, and as we mentioned in the previous section, we will use bi-directional LSTM in this case. We used AG News [22], a benchmark dataset to test our method primarily in the NLP domain. Let us observe the validation loss and accuracy curves generated using the AG News dataset first.
Figure 13: Validation loss curves for the traditional & the D2C method for different numbers of subsets on AG News.

### v2

@@ 3.4. Uncertainty-Aware Global Aggregation and Parameter Update
1:Input: Training data $D$, Validation data $D_{val}$; Num. subsets NN; Epochs $E,E_{global}$; Batch $B$; Learning rate lrlr; Parameters: $\lambda,C_{max},\delta_{e},K,\epsilon_{s}$
2:Output: Central model $M_{c}$ with aggregated weights $\theta_{c}$
3:
4:function DUA-D2C($D,D_{val},N,E,E_{global},B,lr,\lambda,C_{max},\delta_{e},K,\epsilon_{s}$)
@@
5:     Shuffle $D$; Divide $D$ into NN subsets: $\{D_{1},\dots,D_{N}\}$ maintaining class distributions
6:     Initialize central model $M_{c}$ with weights $\theta_{c}$
7:     Set min. inverse entropy: $inv\_e_{min}=1/(\log(K)+\delta_{e})$
8:for $g=1$ to $E_{global}$ do
9:      $W_{central}=0$; $S_{raw}=[\,]$
10:      Let $\Theta^{\prime}_{edge}=[\,]$ be a list to store edge model weights
11:  for $i=1$ to NN do
12:        Initialize edge model $M_{i}$ with $\theta_{i}=\theta_{c}$
13:   for $e=1$ to $E$ do
14:          Train $M_{i}$ on $D_{i}$ (batch $B$, learning rate lrlr)
15:   end for
16:
17:        Compute $a_{i}$ (accuracy of $M_{i}$ on $D_{val}$)
18:        Compute $\epsilon_{i}$ (prediction entropy of $M_{i}$ on $D_{val}$)
19:        $inv\_e_{i}=1/(\epsilon_{i}+\delta_{e})$
20:        $inv\_e_{cap_{i}}=\min(inv\_e_{i},C_{max})$
21:        $u_{i}=(inv\_e_{cap_{i}}-inv\_e_{min})/(C_{max}-inv\_e_{min}+\epsilon_{s})$
22:        $u_{i}=\max(0,\min(1,u_{i}))$ {Clamp to [0,1][0,1]}
23:        $s_{i}=\lambda\cdot a_{i}+(1-\lambda)\cdot u_{i}$
24:        Append $s_{i}$ to $S_{raw}$
@@
25:        Obtain updated weights $\theta^{\prime}_{i}$ from $M_{i}$; Append $\theta^{\prime}_{i}$ to $\Theta^{\prime}_{edge}$
26:  end for
27:
28:      $Sum\_S_{raw}=\sum_{j=1}^{N}S_{raw}[j]$
29:      Let $S_{norm}=[\,]$ be a list for normalized scaling factors
30:  for $i=1$ to NN do
31:        $S_{norm}[i]=S_{raw}[i]/(Sum\_S_{raw}+\epsilon_{s})$
32:  end for
33:
34:  for $i=1$ to NN do
35:        $\theta_{scaled_{i}}=S_{norm}[i]\cdot\Theta^{\prime}_{edge}[i]$
36:        $W_{central}=W_{central}+\theta_{scaled_{i}}$
37:  end for
38:
39:      $\theta_{c}=W_{central}$ {Update central model weights}
40:      Update $M_{c}$ with $\theta_{c}$
41:end for
42:
43:return central model $M_{c}$
44:end function
3.5. Global Hyperparameter Tuning for DUA-D2C
Tuning the DUA-D2C model involves a structured approach. First, the base architecture for the edge models (and thus the central model) and its associated learning hyperparameters (e.g., learning rate, batch size for local training) should be reasonably optimized. This can be done by training a single instance of this architecture on a representative portion of the training data or the entire training set initially, using the validation set for guidance.
Once a suitable base model architecture is established, the DUA-D2C specific global hyperparameters require tuning. These include:
  • Number of Subsets ($N$): The training data is divided into $N$ shards. The optimal $N$ depends on dataset size, complexity, and class distribution.
  • Number of Local Epochs ($E$): The number of epochs each edge model trains on its subset before aggregation. A small number (e.g., $1\sim 3$) is often preferred, as edge models can quickly overfit their reduced-size training subset.
  • Tradeoff Parameter ($\lambda$): This parameter balances the contribution of validation accuracy ($a_{i}$) versus scaled confidence ($u_{i}$) in determining the edge model weights. Its range is typically $[0,1]$. In our experiments, we found the range $0.5\sim 0.8$ to be effective.
  • Max Inverse Entropy Cap ($C_{max}$): This caps the raw inverse entropy value before scaling, preventing extremely confident (but potentially incorrect) models from dominating. This value may be dataset-dependent.
  • Other parameters like $\delta_{e}$ (entropy smoothing) and $\epsilon_{s}$ (scaling stability) are typically small constants and less frequently tuned, but their values should be noted.
The tuning process can proceed by first determining a sensible range for $N$ and $E$. One might start by setting $E$ to a small value (e.g., 1 or 2) and then varying $N$, observing performance metrics like validation accuracy/loss, F1 score, etc., of the final central model. Once a reasonable $N$ is found, $E$ can be tuned. Subsequently, $\lambda$ can be optimized by observing the performance metrics too. $C_{max}$ can be set to a reasonable value solely by observing the typical range of inverse entropy for a particular dataset. The objective is to find the combination that yields the best generalization. The most influential hyperparameter among these four is the number of subsets $N$, as we can see from the mathematical intuition.
4. Experimental Setup
4.1. Datasets
@@
We selected a range of benchmark datasets from three different domains to test and validate our proposed method. For image classification tasks, we employed three widely used benchmark datasets: CIFAR-10 (krizhevsky2014cifar, 15), MNIST (lecun1998gradient, 16), and Fashion MNIST (xiao2017fashion, 29). CIFAR-10 provides a challenging image classification task with 60,000 $32\times 32$ RGB images across 10 categories. MNIST offers a simpler task with 70,000 $28\times 28$ grayscale images of handwritten digits. Fashion MNIST presents a more complex variation with 70,000 $28\times 28$ grayscale images of fashion items. To address class imbalance and the overfitting challenge in facial expression recognition, we utilized the FER-2013 (zahara2020facial, 31) dataset, which comprises over 35,000 grayscale images of seven basic emotions. The diverse and noisy nature of this internet-sourced dataset makes it particularly suitable for evaluating generalization performance. For audio-based tasks, we created a comprehensive dataset by combining TESS (dupuis2010toronto, 23), CREMA-D (cao2014crema, 4), and RAVDESS (livingstone2018ryerson, 18). This dataset encompasses 14 emotion classes, offering a diverse range of emotional expressions and recording conditions. By merging these datasets, we aim to improve model generalization by mitigating biases and limitations inherent in individual datasets. To test our method on textual data, we employed the AG News (zhang2015character, 33) dataset, which consists of approximately 127,600 news articles categorized into four balanced classes: World, Sports, Business, and Science/Technology. This dataset provides a large-scale, real-world benchmark for evaluating text classification models, particularly in the context of Big Data.
4.2. Experimental Details
@@
1. Image Classification: At first, all the images are reshaped and normalized. We applied one-hot encoding to the labels to use cross-entropy loss later. We split the training data into training and validation sets using Scikit-Learn. We then divided the training set into multiple subsets. Before constructing the model, processing each training subset into tensors, and batching them was our last step. Each of our edge models was comprised of 4 Convolutional (32 to 256 kernels progressively) and MaxPooling layer (2X2) blocks, followed by one fully connected hidden layer (128 neurons). Each block contained two identical convolutional layers and one pooling layer.
@@
In all cases, we used a 90-10 Training–Validation split, Adam optimizer, and applied dropouts and batch normalization layers. This is important to test for DUA-D2C’s effectiveness when used on top of other methods that address overfitting. We used the predefined test sets for testing.
5. Results and Discussions
We will use loss and accuracy curves, Accuracy, F1 Score, MCC, Cohen’s Cappa, AUC-ROC, and log loss to analyze the results and evaluate our method. We also varied the two new global hyperparameters we mentioned in the previous section: the Number of Subsets of the Training Set and the Number of Epochs before each round of Central Averaging. To refer to them concisely, we will use the variables $\boldsymbol{N}$ and $\boldsymbol{E}$ respectively in this section. As for the Tradeoff Parameter ($\lambda$), we used a value of 0.7 in each case for the sake of consistency, and due to the fact that we found this value to be a good one.
5.1. Visualizing Decision Boundary
At first, we will visualize the change in decision boundaries due to applying our proposed method and analyze the findings. We used a synthetic binary classification dataset to simulate a simple decision boundary. This dataset was created using the Scikit-Learn library. Then the 240-sample large dataset was divided into a training and a test set using a 75-25 split. The neural network architecture we used for these experiments comprised three hidden layers having 100 neurons each. No regularization or data augmentation was used for decision boundary visualisation, to observe the effect of the Divide2Conquer approach explicitly. For decision boundary analysis, we kept things simple and used the basic D2C approach without performance-based aggregation.
At first, we adopted the centralized, traditional approach and fed the training dataset to the sole model. After training, we got the decision boundary shown in Figure LABEL:fig:_n1-dbl.
@@ 5.1. Visualizing Decision Boundary
Figure 3. Decision boundary visualization with the Traditional approach vs. Divide2Conquer (D2C).
Figure 4. Decision Boundary from the edge models trained on each subset after dividing the training data into 3 subsets using the proposed method.
As we can see, the decision boundary for the traditional approach is highly complex and non-linear here. The model creates intricate regions to distinguish between the two classes. The boundary closely follows individual points, resulting in a convoluted, jagged shape. The complexity of the decision boundary suggests that the model is overfitting to the training data. It is trying too hard to separate every point, including noise, rather than focusing on general patterns in the data. Evidence of overfitting is seen in how the model creates small pockets of blue in the red region and vice versa. This indicates the model has likely memorized the data rather than learning generalizable patterns. This means that this approach is likely to produce excellent results on training data. Still, it is likely to struggle with new, unseen data due to its sensitivity to specific details in the training set. The highly irregular boundary reflects poor generalization ability.
Then, to see the effect of D2C on decision boundaries, we divided our training set into multiple subsets and trained them in parallel using identical Edge Models. Figure 4 shows the decision boundaries we got from each edge model.
@@
To verify, the trained weights from each subset were averaged to get the final weights for the Central Model. The resulting decision boundary is shown in Figure LABEL:fig:_n3-db.
We can see that the decision boundary from the central model of our method is much simpler, dividing the feature space almost diagonally. This suggests a less flexible, more generalized model due to averaging the weights from multiple subsets. This boundary is smooth and does not follow the exact positions of individual data points as closely. The model is more generalized, focusing on the overall trend in the data rather than trying to accommodate every single datapoint, exhibiting much less overfitting compared to what we saw in the traditional approach using the same model architecture. By averaging the weights from different subsets, the model smooths out the noise in the data and prevents overfitting to specific training samples. An individual noisy sample does not affect all the edge models in D2C, and its impact is diluted through averaging. This leads the central model to create a decision boundary based on the more common patterns present in all the subsets, ensuring better generalization. Further proof of this generalization is seen in the training and test accuracies for the traditional approach and D2C using different numbers of subsets. We summarize these accuracies in Table 1.
Table 1. Performance of the traditional approach vs different values of NN Using the proposed method.
@@
Our observations prove that models using the D2C approach perform better on unseen data, as they avoid overfitting by keeping the decision boundary simple and focusing on capturing the broader, more consistent patterns in the data. So, D2C can be seen as yet another technique to reduce overfitting. However, several other techniques address overfitting. So, whether using our proposed method on top of other regularization and augmentation techniques (such as dropouts) improves the performance of a model is the key question at the application level. In the following subsections, we will investigate this using the architectures (incorporating regularization methods like dropouts in the edge models) we described in the previous section on several datasets from different domains. For all these experiments, we used the refined DUA-D2C approach.
5.2. Analyzing Loss & Accuracy Curves
@@
We will analyze the results obtained from our experiments from the 3 different domains. The loss curve is a definitive indicator of overfitting. As we saw in Eq. (6), The validation/test loss truly represents the true error of the model. So we will monitor the validation loss curves from the different experiments we performed to compare and analyze if DUA-D2C improves the generalizing performance, in the form of minimizing the validation loss as the training progresses.
@@
Image Classification: Let us first take a look at the loss curves generated using the CIFAR-10, Fashion MNIST, MNIST, and FER2013 datasets. For now, we will keep the number of epochs before each round of central aggregation, $\boldsymbol{E}$, equal to $1$. It will ensure a proper comparison between the traditional method and our proposed method while varying the number of subsets, $\boldsymbol{N}$. In our experiments, how much we varied $N$ depended on the performance (Loss curve, Accuracy) trend as we increased $N$, the number of subsets. In the case of all the datasets, we stopped varying $N$ whenever we observed a definitive and consistent performance drop with increasing $\boldsymbol{N}$. For example, in the case of CIFAR-10, we varied $N$ in the range of $1$ to $9$.
@@
Figure 5. Validation loss curves for the traditional & the DUA-D2C method for different numbers of subsets, N, on CIFAR-10.
@@
Figure 6. Validation loss curves for the traditional & the DUA-D2C method for different numbers of subsets, N, on Fashion MNIST.
@@
Figure 7. Validation loss curves for the traditional & the DUA-D2C method for different numbers of subsets, N, on MNIST.
@@
Figure 8. Validation loss curves for the traditional & the DUA-D2C method for different numbers of subsets, N, on FER2013.
@@
As we can see in Figures 5, 6, and 8, a key indicator of overfitting, an increasing validation loss after initial training, is notably delayed and mitigated when using our proposed DUA-D2C method compared to traditional implementations. In machine learning, overfitting is often marked by a model’s increasing validation loss after a certain number of training epochs, indicating that the model has started to capture noise and specific variations in the training set that do not generalize. A model that reaches this overfitting point later, as DUA-D2C does, suggests it remains in a generalization phase for longer. We can also see that, in all these cases, dividing the training data into more subsets with DUA-D2C delays and slows down the abrupt increase in validation loss. For example, on CIFAR-10, the traditional approach (no subsets) reached its minimum loss of $0.536$ at around $16$ epochs, then increased to $0.77$ by 100 epochs. With DUA-D2C ($E=1$), for $N=3$, the minimum loss was $0.462$ around $21$ epochs, rising to $0.64$. When $N=9$, the minimum was $0.47$ around 46 epochs, reaching $0.585$. So, clearly, the convergence to the overfitting point becomes slower, and the abrupt increase of validation loss is minimized as we divide the training set into more subsets, which indicates reduced overfitting with DUA-D2C.
Also, DUA-D2C manages to maintain a lower validation loss in general in all cases, including MNIST, as we can see from Figure 7. The minimum value of loss achieved also tells a similar story: we can observe the minimum validation loss for a DUA-D2C variant in each case. When a model maintains a lower validation loss, it indicates better generalization, as it effectively captures true underlying patterns without becoming excessively sensitive to training set noise. This stronger alignment with the data’s general structure rather than its idiosyncrasies means the model is overfitting less. The most striking illustration of this phenomenon is seen with the FER2013 dataset (Figure 8), a task highly prone to overfitting. Here, increasing the number of subsets with DUA-D2C clearly diminishes this overfitting tendency.
This empirical evidence, that employing more subsets helps reduce overfitting, aligns with our theoretical intuition from Eq. (13). As $k$ (number of subsets) increases, the variance of the aggregated model can decrease (approaching $s/k$ in the ideal scenario), directly combating a primary component of overfitting. Thus, the observed minimization of validation loss (our proxy for true error) with DUA-D2C, particularly with a larger $N$, validates our hypothesis and confirms its effectiveness in treating the problem of overfitting.
@@
There is another interesting aspect of the validation loss curve here. As we can see from Figure 7, the validation loss curve generated using the traditional approach deviates much more than the ones generated using the DUA-D2C method. A curve that does not fluctuate much after converging suggests that the model has reached a stable state. This stability is often a sign of robustness, meaning the model is not overly sensitive to slight variations in the data or training process. Dividing the dataset into more and more subsets results in smaller and smaller deviations in the validation loss curve. This was evident in previous loss curves too, to a lesser extent. This proves that dividing the training set into multiple subsets not only improves accuracy and loss but also brings stability and consistency to the loss/accuracy curves.